OpenStudy (mendicant_bias):

@dan815 45 minutes before my exam, trying to crunch in some solutions prior.

OpenStudy (mendicant_bias):

\[\mathcal{L^{-1}}\left\{\vphantom{}\frac{3s-1}{s^2-s-6}\right\}\]

OpenStudy (mendicant_bias):

@mathstudent55 I'm thinking of splitting them up into two separate Laplaces by the numerator, but I'm not sure how to move forward from there.

OpenStudy (mendicant_bias):

@dumbcow

OpenStudy (mendicant_bias):

\[ \frac{3s}{s(s-1)-6} -\frac{1}{s(s-1)-6}\]

OpenStudy (anonymous):

Partial fractions: \[\frac{3s-1}{s^2-s-6}=\frac{A}{s-3}+\frac{B}{s+2}\]

OpenStudy (mendicant_bias):

Heyoo, I think I passed, lol, looks like I won't have to deal with this for a little.